# Optimal Scaling Needs Optimal Norm

### v1

@@ 5 Conclusion and Discussion
In this work, we demonstrate that the operator norm of the output layer is a powerful measure that guides joint optimal scaling across both model and dataset dimensions. Informally, one may view our results as:
@@
  • Which norm is exactly optimal? We paid most of our attention to $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$, but are the observed phenomena really specific to this one only? And to the Scion optimizer only?
@@
We don’t yet have answers to those questions, but we believe our study scratches the surface of exciting phenomena that remain to be fully understood.
@@ A.10 Ablation of depth transfer techniques
A.11 Ablations on Fig. 2(a)
A.11.1 Momentum & learning rate decay
@@ A.11.2 Norm choice
A.12 Learning rate layout for additional batch sizes and horizons
@@
A.13 Individual norm scans and fits

### v2

@@ 5 Conclusion and Discussion
In this work, we demonstrate that the operator norm of the output layer is a powerful measure that guides joint optimal scaling across both model and dataset dimensions. Informally, we show:
@@
  • Which norm is exactly optimal? We paid most of our attention to $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$, but are the observed phenomena really specific to this one only?
@@
We don’t yet have answers to those questions, but we believe our study scratches the surface of exciting phenomena to be further understood.
@@ A.10 Ablation of depth transfer techniques
A.11 Ablation with FineWeb-2 dataset
Figure 9: Same as Fig. 2(a) but using the FineWeb-2 dataset: (a) Thai partition, (b) Russian partition. We note that while for the Russian language the three horizon curves are nested inside of each other and share the same optimal norm, for the Thai partition the first horizon is off. This may be due to being an early phase of training or statistical fluctuation. For completeness, we provide the individual learning rate scans and fits used to produce this plot in Fig. 10.
Figure 10: Individual $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$ norm scans for various batch sizes $B$ (columns) across various horizons $D$ (rows), for the FineWeb-2 dataset: (a) Thai partition, (b) Russian partition.
A.12 Ablation with Adam optimizer
Figure 11: Same as Fig. 2(a) but using the Adam optimizer (no weight decay, warmup for $2^{29}\approx 537$M tokens across all batch sizes, followed by constant learning rate schedule with linear decay to 0 for 20% of the total schedule): (a) $\beta_{1}=\beta_{2}=0$, (b) $\beta_{1}=0.9,\,\beta_{2}=0.95$. For the no-momentum version (a) we observe norm transfer at the same optimal norm value as our main experiment with Scion (Fig. 2(a)). For the version with momentum (b), norm transfer is also present (with reduced sensitivity to batch size choice similarly to Scion), but notably at a higher optimal norm value.
A.13 Residuals-only data scaling
Figure 12: Same as Fig. 2(a) but leaving in the proxy model only residuals normalization layers (RMSNorm without trainable parameters normalizing inputs to every attention and MLP blocks).
A.14 Normalization layers
Figure 13: Learning rate scan for various layer normalization strategies for the proxy model and the Scion optimizer: (a) without momentum $\alpha=1.0$, (b) with momentum $\alpha=0.1$. All runs share the same batch size $B=256$ [sequences] and the same learning rate schedule: no warmup, constant phase ($2^{35}$ tokens), linear decay to 0 for 20% of the total horizon ($0.25\times 2^{35}$ tokens). Learning rate plotted on the X axis corresponds to the learning rate of the constant phase. Markers which are not present on the plot mean that the training diverged.
Notations in the legend corresponds to normalising 1) for the attention block (A) QKV: outputs of the QKV projection matrices, O: inputs to the output projection matrix; 2) MLP: inputs to the last linear layer; 3) residuals (R): inputs to both attention and MLP blocks; 4) Output layer (O): inputs to the final layer projecting the model dimension onto the vocabulary. Additional legend block shows the final norm ($\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$) value and training loss for the top-3 optimal learning rate runs within each normalization strategy.
A.15 Ablations on Fig. 2(a)
A.15.1 Momentum & learning rate decay
@@ A.15.2 Norm choice
A.16 Learning rate layout for additional batch sizes and horizons
@@
A.17 Individual norm scans and fits